Deandre rented a truck for one day. There was a base fee of $16.95 , and there was an additional charge of 96 cents for each mile driven. Deandre had to pay $273.27when he returned the truck. For how many miles did he drive the truck?
Answer by checkley79(3341)

16.95+.96X=273.27
.96X=273.27-16.95
.96X=256.32
X=256.32/.96
X=267 MILES WERE DRIVEN.
PROOF:
16.95+.96*267=273.27
16.95+256.32=273.27
273.27=273.27